A pediatrician is concerned about childhood wrist and forearm fractures affecting fine motor skills later in life. She has approached you to determine whether young adults with a childhood history of wrist/forearm bone fractures are more likely to have fine motor problems. You collected data from 289 patients (aged 18 to 24 years) from 20 clinics in Georgia and the results are summarized below:

No fine motor problems Fine motor problems Total
No history of bone fracture 163 9 172
History of bone fracture 103 14 177
Total 266 23 289
  1. Among those with no history of bone fracture, the prevalence (risk) of fine motor problems, rounded to 0.01, is:
  2. Among those with no history of bone fracture, the odds of fine motor problems, calculated from the table, rounded to 0.01, is:
  3. Among those with a history of bone fracture, the prevalence (risk) of fine motor problems, rounded to 0.01, is:
  4. Among those with a history of bone fracture, the odds of fine motor problems, rounded to 0.01, is:
  5. Based on the scenario described above, to describe the association between wrist/forearm fracture and fine motor problems, you calculate the odds ratio (calculated from the table, not the rounded answers above):
  6. To be comprehensive, the pediatrician asks you to also calculate the risk ratio (remember, the outcome is the presence of fine motor problems and the exposure of interest is a history of bone fracture). This risk ratio (calculated from the table, not the rounded answers above) is:
  7. You are feeling good and now want to flex your epidemiology muscles a bit more. You decide to calculate the risk difference (calculated from the table, not the rounded answers above), rounded to the nearest 0.01, and present:
  8. The pediatrician wants to discuss the relationship between the odds ratio and the risk ratio. You start by saying (select one best answer):

a. The risk ratio always overestimates the odds ratio and should not be reported lest it raise undue anxiety

about the risk of fine motor abilities after bone fracture.

b. The risk ratio is slightly lower compared to the odds ratio, but the inference remains the same: those with a

history of wrist/forearm fracture in childhood were more likely to have fine motor problems as young adults.

c. In this case, the odds ratio and the risk ratio are identical, which is an appealing property of the odds ratio.

d. If one assumes that the outcome is rare, it is better to avoid using the odds ratio.
Answer #1
Fine motor problems no fine motor problems
History of bone fracture 14 (a) 103 (b)
No history of bone fracture 9 (c) 163 (d)

Create a table with the given values of collected data

To calculate the risk ratio and odd ratio,

Risk ratio = a/a+b÷ c/ c+d = 14/(14+103)÷9/(9+163)= 1.6

Odd ratio = a/c÷b/d = 14/9÷ 103/163 = 2.46

To discuss about the relationship between odd ratio and risk ratio, the paediatrician can start with,

b ) The risk ratio is slightly lower compared to the odds ratio, but the inference remains the same: those with ai History of wrist/forearm fracture in childhood were more likely to have fine motor problems as young adults.

Usually odd ratio can overestimate and magnify the risk especially when the disease is more common.